# DySK-Attn: A Framework for Efficient, Real-Time Knowledge Updating in Large Language Models via Dynamic Sparse Knowledge Attention

### v2

@@ 5.2 Ablation Studies
(a) Importance of Sparse Attention: When we replace our sparse knowledge attention mechanism with a standard (dense) attention over all facts in the retrieved subgraph $\mathcal{G}_{sub}$ (variant ‘w/o Sparse Attention‘), we observe a significant performance drop of 7.7 points in F1 score. This demonstrates the effectiveness of our sparsity-driven approach. By forcing the model to select only the top-$k$ most relevant facts, we mitigate the noise from less relevant information within the candidate subgraph. This hard selection acts as a powerful inductive bias, compelling the model to be more decisive and focused in its knowledge utilization, a principle that mirrors the importance of attention mechanisms in fine-grained recognition tasks gu2023attention.

### v3

@@ 5.2 Ablation Studies
(a) Importance of Sparse Attention: When we replace our sparse knowledge attention mechanism with a standard (dense) attention over all facts in the retrieved subgraph $\mathcal{G}_{sub}$ (variant ‘w/o Sparse Attention‘), we observe a significant performance drop of 7.7 points in F1 score. This demonstrates the effectiveness of our sparsity-driven approach. By forcing the model to select only the top-$k$ most relevant facts, we mitigate the noise from less relevant information within the candidate subgraph. This hard selection acts as a powerful inductive bias, compelling the model to be more decisive and focused in its knowledge utilization, a principle that mirrors the importance of attention mechanisms in fine-grained recognition tasks .